# OmniHuman-1: Rethinking the Scaling-Up of One-Stage Conditioned Human Animation Models

### v2

@@ 3.4 Inference Strategies
Figure 5: Subjective comparison of different training ratios for audio conditions.
Methods Identity Consistency Lip-sync Accuracy Visual Quality Action Diversity Overall
@@
[…]
4 Experiments
4.1 Implementation Details
Dataset. By filtering based on aesthetics, image quality, motion amplitude, etc. (common criteria for video generation), we obtained 18.7K hours of human-related data for training. Of this, 13% was selected using lipsync and pose visibility criteria, enabling audio and pose modalities. During training, the data composition was adjusted to fit the omni-condition training strategy. For testing, we conduct the evaluation following the portrait animation method Loopy [26] and the half-body animation method CyberHost [34]. We randomly sampled 100 videos from public portrait datasets, including CelebV-HQ [83] (a diverse dataset with mixed scenes) and RAVDESS [28] (an indoor dataset including speech and song) as the testset for portrait animation. For half-body animation, we used CyberHost’s test set, which includes a total of 269 body videos with 119 identities, encompassing different races, ages, genders, and initial poses.
Figure 6: Ablation study on different audio condition ratios. The models are trained with different audio ratios (top: 10%, middle: 50%, bottom: 100%) and tested in an audio-driven setting with the same input image and audio.
Figure 7: Ablation study on different pose condition ratios. The models are trained with different pose ratios (top: 20%, middle: 50%, bottom: 80%) and tested in an audio-driven setting with the same input image and audio.
Figure 8: Ablation study on different pose condition ratios. The models are trained with different pose ratios (top: 20%, middle: 50%, bottom: 80%) and tested in an audio-driven setting with the same input image and audio.
Figure 9: Ablation study on reference condition ratios. Comparisons of visualization results for 30s videos at different reference ratios.
4.2 Comparisons with Existing Methods
As shown in the Table 5 and 5, overall, OmniHuman demonstrates superior performance compared to leading specialized models in both portrait and body animation tasks using a single model. For audio-driven animation, the generated results cannot be identical to the original video, especially when the reference image contains only a head. The model’s varying preferences for motion styles across different scenarios complicate performance measurement using a single metric. By averaging the metrics across the dataset, OmniHuman achieves the best results across all evaluated metrics, reflecting its overall effectiveness. Additionally, OmniHuman excels across almost all metrics in specific datasets. Notably, existing methods use a single model for specific body proportions (portrait, half-body) with fixed input sizes and ratios. In contrast, OmniHuman supports various input sizes, ratios and body proportions with a single model, achieving satisfactory results. This advantage stems from its omni-conditions training, which learns from a large scale of diverse content and varying sizes during mixed data training.
@@ 4.2 Comparisons with Existing Methods
4.3 Ablation Studies on Omni-Conditions Training
Here, we primarily analyze and explain principles 1 and 2 of the omni-condition training in OmniHuman. For the first principle, we compare training using only data that meets the requirements for audio and pose animation (i.e., 100% audio training ratio) with training data for weaker conditions (i.e., text). Our experimental results demonstrate that the ratio of these two data parts significantly affects the final performance. From the visualizations in Figure 6, it is evident that a high proportion of audio condition-specific data training reduces dynamic range and can cause failures with complex input images. Including weaker condition data at a 50% ratio yields satisfactory results (e.g., accurate lip-syncing and natural motion). However, excessive weaker condition data can hinder training, resulting in poorer correlation with the audio. We also conducted a subjective evaluation to determine the optimal mix of these two data types during training. Specifically, we conducted a blind evaluation with 20 subjects who compared the samples across various dimensions to select the most satisfactory one, with an option for abstention. In total, 50 samples depicting diverse scenarios were evaluated. The results in Table 5 were consistent with the conclusions drawn from the visualizations.
The second principle can also be simultaneously validated with the principle 1 experiment, but we additionally conduct another experiment using different ratios of pose conditions to study the effects of pose condition ratios. Visual comparisons are presented in Figure 7 and  8. When the model is trained with a low pose condition ratio and tested with only audio conditions, the model tends to generate intense, frequent co-speech gestures, as is proven by the motion blur effects in the top row of Figure 8 and the incorrect fingers in the top row of Figure 7. On the other hand, if we train the model with a high pose ratio, the model tends to rely on the pose condition to determine the human poses in the generated video. Consequently, given the input audio as the only driving signal, the generated results typically maintain a similar pose, as shown in the bottom rows of Figure 7 and  8. Therefore, we set the pose ratio to 50% as our final training configuration.
Apart from analyzing the training ratios of new driving modalities in Stage 2 and Stage 3, the training ratio of the appearance condtion is equally important. We investigated the impact of reference image ratios on the generation of 30-second videos through two experiments: (1) setting the reference image ratio to 70%, lower than the text injection ratio but higher than audio; (2) setting the reference image ratio to 30%, lower than the injection ratios for both audio and text. The comparative results are shown in Figure 9, revealing that a lower reference ratio leads to more pronounced error accumulation, characterized by increased noise and color shifts in the background, degrading performance. In contrast, a higher reference ratio ensures better alignment of the generated output with the quality and details of the original image. This can be explained by the fact that when the reference image training ratio is lower than that of audio, the audio dominates the video generation, making it difficult to maintain the ID information from the reference image.
4.4 Extended Visual Results
In the Figure 10 and Figure 11, we present more visual results to demonstrate OmniHuman’s powerful capabilities in human animation, which are difficult to capture through metrics and comparisons with existing methods. OmniHuman is compatible with diverse input images and maintains the motion style of the input, such as preserving the characteristic mouth movements in anime. OmniHuman also excels in object interaction, generating videos of singing while playing different musical instruments and natural gestures while holding objects. Due to its compatibility with pose conditions during training, OmniHuman can perform pose-driven video generation or a combination of pose and audio-driven generation. More video samples can be seen on our project page (highly recommended).
Figure 11: The videos generated by OmniHuman based on input audio and images. These demonstrates OmniHuman’s compatibility with various environments, objects, and camera angles, producing satisfactory results.
@@ 5 Conclusion
We propose OmniHuman, an end-to-end multimodality-conditioned human video generation framework that generates human videos based on a single image and motion signals (e.g., audio, video, or both). OmniHuman employs a mixed data training strategy with multimodality motion conditioning, leveraging the scalability of mixed data to overcome the scarcity of high-quality data faced by previous methods. It significantly outperforms existing approaches, producing highly realistic human videos from weak signals, especially audio. OmniHuman supports images of any aspect ratio (portraits, half-body, or full-body) delivering lifelike, high-quality results across various scenarios.

### v3

@@ 4.1 Implementation Details
4.2 Ablation Studies on Omni-Conditions Training
Principles 1. Here, we first analyze and explain Principle 1 of the omni-condition training. This principle guides us to introduce more training data using the weaker text condition compared to the commonly used audio and pose conditions in human animation tasks. In the upper part of Table 1, we present the results of using varying proportions of text-conditioned data (T-DATA) on the results. It can be seen that more text-conditioned data improves core metrics such as FVD, FID, lipsync accuracy, and the richness and quality of gesture generation. We observed that when the amount of text-conditioned data is insufficient, the lipsync metric declines. This may be because a small amount of new data disrupts audio learning rather than enhancing text condition capability. As the data increases, the model leverages the shared backbone to learn better joint modality modeling, resulting in improved generated results. We also found that with continued training and increased data, the model gradually generates results that follow the quality of the input images rather than the distribution quality of the training set. Interestingly, while the image quality score may appear to decrease, the model actually learns to adhere to the distribution of the input images rather than the higher quality, limited audio-conditioned training data distribution. The introduction of text-conditioned data not only improves metrics but also significantly enhances the model’s generalization to different input types.
Principles 2. This principle guides us on how to appropriately handle the relationships between different modality conditions during training, ensuring they complement each other to achieve satisfactory results. Since the sequential introduction of text and image in training is a common practice in the field of video generation, we follow this setup. However, the order of introducing pose and audio during training differs. Previous works [60, 26, 34, 38] have introduced pose or pose-related features alongside audio features in the training process to stabilize it. By adhering to Principle 1, OmniHuman leverages a large amount of text-conditioned data to learn general video generation capabilities. We found that introducing pose too early in the training process can actually harm the model’s ability to generate high-quality motion. For convenience, we use I, T, A, and P to represent the modalities of reference image, text, audio, and pose, respectively and their training order. We use IA to denote our pure audio-driven model. In the lower part of Table 1, we observe that training with pose first (IPA) leads to a comprehensive decline in the model’s final output quality across multiple dimensions, affecting the upper limit of the model’s performance. Conversely, introducing pose later in the training process (IAP) does not harm the model’s results, achieving comparable metrics and enabling the model to support both independent and mixed audio and pose driving. As shown in Figure 7, we also found that the hybrid-driven training model IAP decouple hand movements from the associated audio. This mitigates the issue of excessively exaggerated hand movements in the generated videos, significantly enhancing the naturalness of human dynamics.
Based on Principle 2, we maintain a lower training proportion for pose compared to audio (A>P, A=50% and P=25%), by adjusting the condition drop rate. The results in Table 1 show that using the opposite training strategy (A<P, A=50% and P=75%) leads to degraded generation quality. We also validated the effects of different training proportions of audio compared to text in Supplementary Section 5. Specifically, we kept the text-condition ratio T=90% constant and experimented with A=10%, A=50%, and A=90%. According to the quantitative and qualitative results, the A=50% ratio yields a more balanced performance, showing satisfactory generalization and lip-sync accuracy.
Effects of omni-condition training. In addition to our investigation into training principles, we also provide a comparison with the results of pretrained image-to-video model, as shown in Figure 5. The proposed omni-condition training allows our model to leverage the relationships between mixed conditions during inference. This significantly enhances the generative capabilities of pretrained I2V models, improving the quality of local human features such as gestures, as well as the overall realism and naturalness of movements in the generated videos. These results demonstrate that the proposed methods are crucial for improving generative performance.
4.3 Comparisons with Existing Methods
In this section, we provide a comparison between OmniHuman and existing methods, including audio-driven methods and pose-driven ones.
Comparison with audio-driven methods. As shown in Tables 4 and 4, OmniHuman demonstrates superior performance compared to leading specialized models in both portrait and body animation tasks using a single model. For audio-driven animation, the generated results cannot be identical to the original video, especially when the reference image contains only a head. The model’s varying preferences for motion styles across different scenarios complicate performance measurement using a single metric. OmniHuman excels across almost all metrics in different datasets and tasks. Notably, existing methods target specific body proportions (portrait, half-body) with fixed input sizes and ratios. In contrast, OmniHuman supports various input sizes, ratios, and body proportions with a single model, achieving satisfactory and realistic results. More than just performance superiority, as far as we know, OmniHuman is the first solution capable of this, which further underscores the effectiveness of the proposed omni-condition training strategy.
Comparison with pose-driven methods. In addition to audio-driven animation, we also provide a comparison of OmniHuman with recent pose driving methods. It is worth noting that these comparison methods are all specifically trained for pose-driven tasks, while OmniHuman, being a unified model can handle different tasks including both audio-driven and pose-driven human video generation. The results in Table 2 indicate that OmniHuman can also perform well on pose-only driving tasks with satisfactory visual quality and driving accuracy.
Table 2: Comparison with video-driven body animation methods.
@@ 4.3 Comparisons with Existing Methods
DisCo [62] 3.707 2.396 57.12 64.52 9.313
AnimateAnyone [25] 3.843 2.718 26.87 37.67 5.747
MimicMotion [76] 3.977 2.842 23.43 22.97 8.536
CyberHost [34] 4.087 2.967 20.04 7.7178 3.123
OmniHuman-1 4.111 2.986 19.504 7.3184 2.136
Figure 5: Visual comparison with the pretrained image-to-video model. Blue circle denotes the I2V model and green circle denotes the OmniHuman.
@@
Figure 7: Visual analysis of the hybrid-driven training model. The gradient curves visualize the hand motion trajectories in the video. Blue circle denotes IA model results and green circle denotes IAP model results.
4.4 Extended Visual Results and Applications
In Figure 6, we showcase OmniHuman’s robust capabilities in human animation, which are challenging to quantify through metrics alone. OmniHuman is compatible with diverse input images and maintains the motion style of the input, such as preserving characteristic mouth movements in anime. Moreover, it excels in object interaction, generating videos of singing while playing instruments and natural gestures while holding objects. By adjusting the regions affected by CFG, OmniHuman can specify the speaker in the video, greatly expanding its application value. This versatility stems from OmniHuman’s incorporation of various conditions and data during training, enabling it to learn general human animation through data scaling, unlike existing methods limited by narrowly filtered training data.
@@ 5 Conclusion
We propose OmniHuman, an end-to-end multimodality-conditioned human video generation framework that generates human videos based on a single image and motion signals (e.g., audio, pose, or both). OmniHuman employs a mixed data training strategy with multimodality motion conditioning, leveraging the scalability of mixed data to overcome the scarcity of high-quality data faced by previous methods. It outperforms existing approaches, producing highly realistic human videos from weak signals, especially audio. OmniHuman supports images of any aspect ratio (portraits, half-body, or full-body), delivering lifelike, high-quality results across various scenarios.
Limitation
OmniHuman mitigates overfitting to audio variations through mixed-conditions training, especially by incorporating pose conditions. However, due to the weak correlation between audio and motion, uncoordinated or overly expressive movements still occur and object interactions can sometimes appear unrealistic. These stem from insufficient training data because when input images differ significantly from the training distribution, the model tends to generate unnatural results. To maintain synthesis stability, a relatively high CFG scale is needed, which can lead to some degree of overfitting. Future work will focus on adding richer motion conditions such as styles, intensities, and intentions to improve the naturalness of the generated motions.